# Improving Code LLM Robustness to Prompt Perturbations via Layer-Aware Model Editing

### v1

@@ Abstract.
copyright: nonecopyright: noneconference: ; ;
@@ 1. Introduction
  • We propose a new task formulation: improving the robustness of LLMs to natural language prompt perturbations via internal model editing.
  • We introduce CREME, a novel editing framework that uses a single pair of original and perturbed prompts to identify robustness-sensitive layers and update targeted parameters to enhance the robustness of LLMs.
@@ 2.1. Motivating Example
In real-world scenarios, developers interact with code LLMs by providing natural language prompts to generate desired code completions. However, users may inadvertently introduce minor errors into these prompts, such as typographical mistakes, repeated words, or slight phrasing inconsistencies. Although such errors are typically semantically negligible and easily overlooked by human programmers, they can cause LLMs to produce significantly different outputs.
@@ 2.2. Task Definition
[…]
@@ 3.2. Key Layer Localization
Layer-wise Intervention. To ensure a fair comparison and stable intervention, we construct a mini-batch that includes one instance of $P_{ori}$ and multiple instances (e.g., five copies) of $P_{pert}$. The batch is tokenized using left padding, ensuring alignment of token positions across all samples. We input the batch into the model $G$. For each transformer layer $L_{\ell}$, where $\ell\in\{1,2,\ldots,N\}$, we intervene in the forward pass as follows:
@@
This ratio captures how much of the accuracy gap between $P_{ori}$ and $P_{pert}$ is recovered by intervention at layer $L_{\ell}$, normalized to the maximum possible improvement. We define the key layer $L_{\ell^{*}}$ as the one with the highest restoration improvement. If multiple layers achieve the highest restoration improvement, we adopt the approach proposed by Wang et al. (Wang et al., 2024b) and select the layer exhibiting the greatest hidden state discrepancy between $P_{ori}$ and $P_{pert}$ when passed through model $G$.
@@ 3.3. Layer Editing
Following the localization of the most critical layer for the degradation of robustness, we proceed to perform parameter-level editing to enhance the robustness of the model to a specific perturbation type.
Each layer’s MLP block in Equation 3 is a two-layer neural network (Dai et al., 2021; Meng et al., 2022a; Wang et al., 2024b), and the second neural network can be expressed as:
@@ 4.3. Evaluation Metrics
[…]
G-RIR is a metric we propose to quantify the generalization ability of a robustness enhancement strategy. Given a set of $N{+}x$ prompts associated with a specific perturbation type, we use $x$ prompts to update the model (e.g., $x=1$ in our method). We then evaluate the resulting robustness improvement on the remaining $N$ prompts, which exhibit the same type of perturbation but are unseen during the model updating. Formally, G-RIR is defined as:
[…]
@@ 5.1. RQ1: Effectiveness
To evaluate the effectiveness of CREME, we compare our method with several baselines, including two robustness enhancement techniques (i.e., Self-Denoising and LoRA fine-tuning) and two knowledge editing methods (i.e., ROME and DINM). We assess performance on two representative perturbation types (C3 and D2), and report both absolute functional accuracy (Pass@1) and robustness generalization (G-RIR) under the leave-one-in evaluation protocol (see Section 4.4).
[…]
Pass@1 Comparison: CREME exhibits strong potential in enhancing model robustness. As shown in Table 2, it achieves the highest Pass@1 scores in 7 out of 8 settings. In all cases, CREME outperforms the unedited model on perturbed prompts, improving the ability of models to generate functionally correct code despite input perturbations. Notably, the Self-Denoising approach performs poorly in most settings, except for QWenCoder+MBPP, contrasting with its previously reported effectiveness (Agrawal et al., 2025). This discrepancy may be attributed to two factors: ❶ unlike like QwenCoder, CodeLlama is not instruction-tuned, which limits its ability to revise perturbed prompts based on meta-instructions; ❷ prompts in HumanEval average 231.6 characters in length, significantly longer than those in MBPP (88.2 characters), which increases the likelihood of semantic drift during denoising.
[…]
G-RIR Comparison: CREME exhibits strong generalization capabilities for robustness enhancement. As shown in Table 3, it achieves the highest G-RIR scores in nearly all evaluated settings. For instance, CREME achieves a G-RIR of 0.5120 under the Qwen + MBPP + D2 configuration, significantly outperforming all baselines. Self-Denoising is excluded from the G-RIR comparison as it is designed to enhance robustness on a single instance and does not support generalization across tasks within the same perturbation category. Additionally, some methods (e.g., ROME and DINM) yield negative G-RIR values in certain settings, suggesting that editing with a single standard answer may lead to overfitting and reduced robustness on the same perturbation.
@@ 5.2. RQ2: Generality
To evaluate the generalization capability of our method across different categories of prompt perturbations, we calculate the G-RIR for each perturbation type listed in Table 4.
Table 4 reports the G-RIR scores for 20 perturbation types, including additions (A1–A3), deletions (D1–D4), edits (E1–E6), swaps (S1–S2), paraphrases (P1–P2), and co-occurring categories (C1–C3). Our key findings are summarized as follows:
@@
Answer to RQ2: Our method demonstrates strong generalization across various categories of perturbations, achieving an average improvement of 30%. It performs better on simpler perturbations (e.g., word-level modifications) than on more complex sentence-level rephrasings.
@@
w/o Layer Localization 0.2296$\downarrow$ 0.2306 0.3003$\downarrow$ 0.2364$\downarrow$ 0.3249$\downarrow$ 0.2594$\downarrow$ 0.3074$\downarrow$ 0.4838$\downarrow$
w/o Early Stopping 0.2593 0.2278$\downarrow$ 0.3305 0.2824 0.3540 0.2634 0.3398 0.4969
@@ 5.3. RQ3: Ablation Study
To investigate the contribution of each component (as described in Section 3) to the effectiveness of CREME, we create the following variants:
@@ 5.4. RQ4: Layer Distribution Analysis
To understand how robustness-related information is localized across the model’s architecture, we analyze the distribution of key layers selected by our causal tracing procedure under different perturbation types and models. Figure 3 presents the heatmaps of key layer frequencies for the CodeLlama-7b and Qwen2.5-Coder-7B models, grouped by six perturbation categories.
[…]
@@ 6.1. Impact of Editing on Model Performance
While CREME effectively enhances robustness against perturbed prompts, it is crucial to evaluate whether the editing process inadvertently degrades model performance on clean data. We conduct a case study by randomly sampling 40 editing tasks from the dataset used in each experimental configuration. To ensure balanced coverage across perturbation types, the sampled tasks include two instances for each perturbation type listed in Table 4. For each task, we apply the CREME editing procedure and then evaluate the edited model on the corresponding original dataset (i.e., HumanEval or MBPP) with clean prompts. Table 6 presents the comparison of pass@1 and pass@10 accuracy before and after editing. The results indicate that the performance of CREME remains stable on clean inputs, with most variations within a $\pm$1margin.
@@ 8. Conclusion
In this paper, we presented CREME, a lightweight model editing framework to enhance the robustness of LLMs against natural language prompt perturbations. By leveraging a pair of original and perturbed prompts, CREME identifies robustness-sensitive layers through causal tracing and applies targeted parameter updates to align their internal representations. We conduct extensive experiments on the HumanEval and MBPP benchmarks using two representative code generation models. The results show that CREME substantially improves functional correctness on perturbed prompts, achieving a 63% increase in Pass@1 accuracy, while maintaining performance on clean inputs within a $\pm$1margin. We further assess generalization and show that CREME restores up to 30% of robustness across perturbation types. Ablation studies confirmed the importance of each component, particularly layer localization. Our analysis of key layer distributions reveals that robustness-sensitive regions are concentrated in the middle-to-deep layers and vary with model architecture, offering actionable insights for future robustness-oriented interventions. CREME provides a practical path toward making LLMs more reliable under real-world prompt variability, without requiring full retraining or architectural modifications. Future work will explore extending CREME to broader task domains and evaluating its effectiveness against more challenging perturbations.

### v2

# CREME: Robustness Enhancement of Code LLMs via Layer-Aware Model Editing

@@ Abstract.
Robustness, Code Generation, Model Editing, Large Language Model
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773111isbn: 979-8-4007-2025-3/26/04ccs: Software and its engineering Software creation and management
@@ 1. Introduction
  • We introduce CREME, a novel editing framework that uses a single pair of original and perturbed prompts to identify robustness-sensitive layers and update targeted parameters to enhance the robustness of LLMs. The replication package of our work is publicly available at (cre, 2025).
@@ 2.1. Motivating Example
In real-world scenarios, developers interact with code LLMs by providing natural language prompts to generate desired code completions. However, users may inadvertently introduce minor errors into prompts, such as typographical mistakes, repeated words, or slight phrasing inconsistencies. Although such errors are semantically negligible and easily overlooked by human programmers, they can cause LLMs to produce significantly different outputs.
@@ 2.2. Task Definition
$P_{ori}\approx P_{pert}\quad\text{(slight perturbations)}\quad\not\Rightarrow\quad$
(1) $G(P_{ori})\approx G(P_{pert})\quad\text{(output equivalence)}$
@@ 3.2. Key Layer Localization
Layer-wise Intervention. To ensure a fair comparison and stable intervention, we construct a mini-batch that includes one instance of $P_{ori}$ and multiple instances (e.g., five copies) of $P_{pert}$. This design allows us to stably compute restoration improvement (as defined in Equation 5) across all layers by averaging over repeated forward passes. The batch is tokenized using left padding, ensuring alignment of token positions across all samples. We input the batch into the model $G$. For each transformer layer $L_{\ell}$, where $\ell\in\{1,2,\ldots,N\}$, we intervene in the forward pass as follows:
@@
This ratio captures how much of the accuracy gap between $P_{ori}$ and $P_{pert}$ is recovered by intervention at layer $L_{\ell}$, normalized to the maximum possible improvement. We define the key layer $L_{\ell^{*}}$ as the one with the highest restoration improvement. If multiple layers achieve the highest restoration improvement, we adopt the approach proposed by Wang et al. (Wang et al., 2024b). We compute the Euclidean distance between the hidden states of each candidate layer when processing $P_{ori}$ and $P_{pert}$, and select the layer with the greatest distance.
@@ 3.3. Layer Editing
Following the localization of the most critical layer for the degradation of robustness, we proceed to perform parameter-level editing to enhance the robustness of the model to a specific perturbation type. Each layer’s MLP block in Equation 3 is a two-layer neural network (Dai et al., 2021; Meng et al., 2022a) and the second neural network can be expressed as:
@@ 4.3. Evaluation Metrics
(11) $\text{Pass@}k=1-\frac{{\binom{n-c}{k}}}{{\binom{n}{k}}},\quad\text{for }k\leq n$
G-RIR is a metric we propose to quantify the generalization ability of a robustness enhancement strategy. Given a set of N+xN{+}x prompts associated with a specific perturbation type, we use $x$ prompts to update the model (e.g., $x=1$ in our method). We then evaluate the resulting robustness improvement on the remaining $N$ prompts, which exhibit the same type of perturbation but are unseen during the model updating. Formally, G-RIR is defined as:
(12) $\text{G-RIR}=\frac{1}{N}\sum_{i=1}^{N}\frac{A^{\text{after}}_{\text{pert},i}-A^{\text{before}}_{\text{pert},i}}{A^{\text{before}}_{\text{orig},i}-A^{\text{before}}_{\text{pert},i}+\epsilon}$
@@ 5.1. RQ1: Effectiveness
To evaluate the effectiveness of CREME, we compare our method with several baselines, including two robustness enhancement techniques (i.e., Self-Denoising and LoRA fine-tuning) and two knowledge editing methods (i.e., ROME and DINM). We assess performance on two representative perturbation types, C3 and D2, which represent different levels of perturbation intensity: D2 introduces a single localized modification, whereas C3 involves multiple dispersed changes. We report both absolute functional accuracy (Pass@1) and robustness generalization (G-RIR) under the leave-one-in evaluation protocol (see Section 4.4).
Σ Pass@1 Comparison: CREME exhibits strong potential in enhancing model robustness. As shown in Table 2, it achieves the highest Pass@1 scores in 7 out of 8 settings. In all cases, CREME outperforms the unedited model on perturbed prompts, improving the ability of models to generate functionally correct code despite input perturbations. Notably, the Self-Denoising approach performs poorly in most settings, except for QWenCoder+MBPP, contrasting with its previously reported effectiveness (Agrawal et al., 2025). This discrepancy may be attributed to two factors: ❶ unlike like QwenCoder, CodeLlama is not instruction-tuned, which limits its ability to revise perturbed prompts based on meta-instructions; ❷ prompts in HumanEval average 231.6 characters in length, significantly longer than those in MBPP (88.2 characters), which increases the likelihood of semantic drift during denoising.
Σ G-RIR Comparison: CREME exhibits strong generalization capabilities for robustness enhancement. As shown in Table 3, it achieves the highest G-RIR scores in nearly all evaluated settings. For instance, CREME achieves a G-RIR of 0.5120 under the Qwen + MBPP + D2 configuration, significantly outperforming all baselines. Self-Denoising is excluded from the G-RIR comparison as it is designed to enhance robustness on a single instance and does not support generalization across tasks within the same perturbation category. Additionally, some methods (e.g., ROME and DINM) yield negative G-RIR values in certain settings, suggesting that editing with a single standard answer may lead to overfitting and reduced robustness on the same perturbation.
@@ 5.2. RQ2: Generality
Table 5. Cross-Category Generalization: G-RIR (%) on Unseen Perturbation Types After Editing on C3 (HumanEval+QwenCoder)
@@
G-RIR 24 19 28 23 21 2 16 9 32 38 25 19 20 3 19 10 15 11 7 23
To evaluate the generalization capability of our method across different categories of prompt perturbations, we calculate the G-RIR for each perturbation type listed in Table 4. Table 4 reports the G-RIR scores for 20 perturbation types, including additions (A1–A3), deletions (D1–D4), edits (E1–E6), swaps (S1–S2), paraphrases (P1–P2), and co-occurring categories (C1–C3). Our key findings are summarized as follows:
@@
Finding 4. To evaluate whether model editing tailored to a specific perturbation category can generalize to others, we conduct an additional cross-category evaluation. Specifically, the model is edited using only C3 type examples under the HumanEval+QwenCoder setting and then tests across all other perturbation types. As shown in Table 5, the model achieves an average G-RIR improvement of 18% on unseen perturbation types, compared to 24% when both editing and evaluation are performed within the same category (as reported in Table 4). These results suggest that CREME does not overfit to a specific perturbation category and can generalize to a broader range of prompt variations.
Answer to RQ2: Our method demonstrates strong generalization across various categories of perturbations, achieving an average improvement of 30%. It performs better on simpler perturbations (e.g., word-level modifications) than on more complex sentence-level rephrasings. Moreover, CREME preserves robustness improvements even when trained on a single perturbation type and evaluated across others, achieving an average cross-type G-RIR of 18%.
@@
[…]
@@ 5.3. RQ3: Ablation Study
To investigate the contribution of each component to the effectiveness of CREME, we create the following variants:
@@ 5.4. RQ4: Layer Distribution Analysis
To understand how robustness-related information is localized across the model’s architecture, we analyze the distribution of key layers selected by our causal tracing procedure under different perturbation types and models. Figure 3 presents the heatmaps of key layer frequencies for the CodeLlama-7b and Qwen2.5-Coder-7B models, grouped by six perturbation categories. The “key layer frequency” refers to how many times each transformer layer is selected as the most robustness-sensitive layer across all editing cases (2 datasets × 20 perturbation types).
@@ 6.1. Impact of Editing on Model Performance
While CREME effectively enhances robustness against perturbed prompts, it is crucial to evaluate whether the editing process inadvertently degrades model performance on clean data. We conduct a case study by randomly sampling 40 editing tasks from the dataset used in each experimental configuration. To ensure balanced coverage across perturbation types, the sampled tasks include two instances for each perturbation type listed in Table 4. For each task, we apply the CREME editing procedure and then evaluate the edited model on the corresponding original dataset (i.e., HumanEval or MBPP) with clean prompts. Table 7 presents the comparison of pass@1 and pass@10 accuracy before and after editing. The results indicate that the performance of CREME remains stable on clean inputs, with most variations within a $\pm$1% margin.
@@
6.2. Costs of CREME
To assess the practical feasibility of CREME, we conducted a cost evaluation on the HumanEval+QwenCoder setting using the C3 perturbation category. Table 8 presents a comparison of CREME with four representative baseline methods in terms of localization, training, and testing time.
The layer localization step of CREME takes an average of 495s (8.3 minutes). While this step introduces additional overhead, it is crucial for identifying the most robustness-sensitive layer. In contrast, the two model editing methods, ROME and DINM, rely on either fixed layers or hidden-state distance, neither of which is robustness-aware. As demonstrated in Table 6, our localization strategy substantially improves robustness over fixed-layer editing, justifying the additional cost. Regarding training time, CREME requires only 5.3s, which is significantly less than DINM (15.8s) and comparable to the others. For inference, CREME’s average test time (3.0s) is competitive and markedly faster than Self-Denoising (29.4s), which iteratively adjusts prompts during generation. Given that models are typically deployed across many downstream tasks after editing, minimizing test time is more critical.
To reduce the overhead of layer localization, we propose a lightweight variant that limits localization to a small set of frequently identified robustness-sensitive layers from Section 5.4 (e.g., layers 28/30 in CodeLlama and 17/26 in QwenCoder), which significantly lowers localization time cost while maintaining effectiveness.
Table 8. Comparison of Time Costs (s) for CREME and Baselines on HumanEval+QwenCoder (C3 perturbation)
[…]
Self-Denoising - - 29.4
LoRa 9.2 2.8
ROME 7.59 3.4
DINM 0.2 15.8 3.9
[…]
6.3. Hyperparameter Sensitivity of $\lambda$
The hyperparameter $\lambda$ in the Equation 8 controls the trade-off between the preservation loss and the alignment loss. To guide practitioners in choosing an appropriate value for $\lambda$, we conducted a sensitivity analysis using the C3 perturbation category under the HumanEval+QwenCoder setting. As shown in Table 9, both G-RIR and Pass@1 scores improve as $\lambda$ increases from 0.01 to 0.1, peaking at $\lambda=0.1$. Beyond this point, performance begins to decline. These findings indicate that $\lambda=0.1$ provides a robust trade-off and can serve as a strong default for future applications of our framework.
Table 9. Sensitivity Analysis of $\lambda$ on HumanEval+QwenCoder (C3 perturbation)
[…]
G-RIR (%) 32.1 33.2 35.0 30.3 32.8 27.6
Pass@1 (%) 41.4 41.9 42.2 40.8 41.2 39.6
@@ 8. Conclusion
In this paper, we presented CREME, a lightweight model editing framework to enhance the robustness of LLMs against natural language prompt perturbations. By leveraging a pair of original and perturbed prompts, CREME identifies robustness-sensitive layers through causal tracing and applies targeted parameter updates to align their internal representations. We conduct extensive experiments on the HumanEval and MBPP benchmarks using two representative code generation models. The results show that CREME substantially improves functional correctness on perturbed prompts, achieving a 63% increase in Pass@1 accuracy, while maintaining performance on clean inputs within a $\pm$1% margin. We further assess generalization and show that CREME restores up to 30% of robustness across perturbation types. Ablation studies confirmed the importance of each component, particularly layer localization. Our analysis of key layer distributions reveals that robustness-sensitive regions are concentrated in the middle-to-deep layers and vary with model architecture, offering actionable insights for future robustness-oriented interventions. CREME provides a practical path toward making LLMs more reliable under real-world prompt variability, without requiring full retraining or architectural modifications. Future work will explore extending CREME to broader task domains and evaluating its effectiveness against more challenging perturbations.
Acknowledgement
This research is supported by the National Key R&D Program of China (No. 2024YFB4506400) and the National Research Foundation, under its Investigatorship Grant (NRF-NRFI08-2022-0002). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore.